Opened 6 years ago

Last modified 5 months ago

#482 new flight processing

EM10/06, flight day 134/2013, Perugia

Reported by: edfi Owned by: benj
Priority: alpha 4 medium Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tipo)

Data location: /users/rsg/arsf/arsf_data/2013/flight_data/italy/EM10_06-2013_134_Perugia

Data arrived from ARSF via network transfer on 21/05/2013

EM10/06 Scientific objective:

EM10/06 Priority: a4m

PI: Ken McCaffrey

Sensors:

Camera (Delivered 9/8/2013)
Eagle (Delivered 9/8/2013)
Hawk (Delivered 9/8/2013)
LiDAR (Delivered 9/8/2013)

Change History (22)

comment:1 Changed 6 years ago by edfi

  • Description modified (diff)

comment:2 Changed 6 years ago by edfi

Changed the prefix of the Eagle and Hawk files to be VNIR and SWIR respectively.

comment:3 Changed 6 years ago by edfi

  • Description modified (diff)

comment:4 Changed 6 years ago by edfi

  • Description modified (diff)

comment:5 Changed 6 years ago by anhi

Finished Nav processing

comment:6 Changed 6 years ago by anhi

  • Description modified (diff)
  • Summary changed from EM10/06, flight day 135/2013, Perugia to EM10/06, flight day 134/2013, Perugia

comment:7 Changed 6 years ago by tipo

Started LiDAR processing

comment:8 Changed 6 years ago by knpa

Beginning RCD processing.

comment:9 Changed 6 years ago by tipo

LiDAR delivery created.

comment:10 Changed 6 years ago by anhi

LiDAR Delivery check:

No problems found.

comment:11 Changed 6 years ago by anhi

SCTs for hyperspectral processing. The navigation was invalid for all lines.

The hdr files GPS data froze at 08:44 so GPS values were obtained from the electronic logsheet and new hdr files were made. This has resulted in much larger SCTs than usual.

Eagle and Hawk line 13 was too small to SCT and there was a note on the logsheet that it is not a proper line so will not be processed and the other lines will be renamed.

Flightline Eagle SCT Hawk SCT
1 2.69 2.53
2 2.88 1.27
3 -31.09 -31.98
4 23.913 23.77
5 25.35 25.04
6 24.08 23.68
7 13.24 12.54
8 14.95 14.05
9 -8.79 -8.32
10 12.20 12.58
11 2.82 2.29
12 3.27 3.34
13 -15.64 -8.63
14 -23.44 -22.05
15 -35.37 -33.21
16 -19.44 -16.47
17 8.52 9.21
18 -33.40 -32.76
19 13.82 14.62
20 -7.74 -7.17
21 35.55 35.44
22 29.99 30.22

comment:12 Changed 6 years ago by anhi

Hyperspectral delivery created

comment:13 Changed 6 years ago by edfi

RCD D/C Complete no problems found

comment:14 Changed 6 years ago by tipo

Delivery checked hyperspectral. No problems found.

comment:15 Changed 6 years ago by tipo

Camera, hyperspectral and LiDAR delivered.

comment:16 Changed 6 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Description modified (diff)
  • Owner set to benj

comment:17 Changed 6 years ago by tipo

Lidar was mistakenly processed as UTM32 intstead of UTM33.
Reprocessing using UTM 33.

comment:18 Changed 6 years ago by tipo

LiDAR reprocessed and new delivery created.

comment:19 Changed 6 years ago by anhi

LiDAR Delivery check:

No problems found, ready for delivery.

comment:20 Changed 6 years ago by tipo

Redelivered LiDAR 21/08/2013

comment:21 Changed 5 years ago by lah

Submitted to NEODC for archiving 16/09/2014.

comment:22 Changed 5 months ago by wja

Archiving
Data is already backed up on CEDA but no JSONs have been created.
Have created JSONs, added them to PostGIS and RSYNCed to CEDA.

Last edited 5 months ago by wja (previous) (diff)
Note: See TracTickets for help on using tickets.